OpenStudy (anonymous):

a toy rocket is launched at an angle of 55 degrees above horizontal with an initial velocity of 150 m/s. what is its initial horizontal velocity and initial vertical velocity. use 1 cm = 10 m/s

OpenStudy (anonymous):

Use 1cm=10 m/s....that makes no sense at all. 1cm=0.01m and is not a unit of velocity. Anyways: \[v_y=\left| \vec{v} \right|\sin \theta=150\sin 55\approx122.9 m/s\]and,\[v_x=\left| \vec{v} \right|\cos \theta=150\cos 55\approx86.0 m/s\]